# Investigating the importance of social vulnerability in opioid-related mortality across the United States

### v2

@@ Discussion
XGBoost model
In this study, XGBoost was used to predict opioid-related mortality rates from 2011 to 2022 based on the previous year’s SVI data. To use XGBoost, several parameters need to be selected, such as the number and depth of trees. To select these parameters, we employed a grid search algorithm with cross-validation, optimizing for the highest predictive accuracy. For training and testing, we used 5-fold cross-validation to randomly split the SVI input data into five folds. In each fold, XGBoost trained on four subsets and predicted on one. This process was repeated across all five folds to generate predictions for every county, ensuring that the model was both trained and tested on the entire SVI dataset.
To describe this model mathematically, fix a year $t\in\{2010,2011,\dots,2021\}$ and let $X^{(t)}=\big{[}x_{i,j}^{(t)}\big{]}\in\mathbb{R}^{3144\times 13}$ be the matrix of SVI features input to the model in year $t$. Each row of matrix $X^{(t)}$ corresponds to a county, while each column corresponds to an SVI variable. In particular, let $x_{i,*}^{(t)}$ denote the row vector containing all the SVI feature data for county $i$, and let $x_{*,j}^{(t)}$ denote the column vector containing all county rates for SVI variable $j$. Letting $Y^{(t+1)}=\big{(}y_{i}^{(t+1)}\big{)}\in\mathbb{R}^{3144}$ be the target vector of mortality rates in year $t+1$, we can define the full dataset $X^{(t)}\times Y^{(t+1)}$ passing through the model to be the set of all counties’ feature vectors together with their corresponding target values:
@@
Starting with the root node, an individual decision tree grows by recursively applying a splitting process to each node $n$ in the tree. Let the data at node $n$ be denoted by $Q_{n}$ with $s_{n}$ samples. Then $s_{n}$ is the number of counties which reached that node, and $Q_{n}\subseteq X^{(t)}\times Y^{(t+1)}$ is the subset of data for those counties. At each node, a potential split $\theta=(j,h_{n})$, consisting of an SVI feature $j$ and a threshold $h_{n}$, partition $Q_{n}$ into two subsets:
$Q_{n}^{left}(\theta)$ =$=\{(x_{i,*}^{(t)},y_{i}^{(t+1)})\,|\,x_{i,j}^{(t)}\leq h_{n}\},$
@@
This process creates child nodes from the subsets $Q_{n}^{left}(\tilde{\theta})$ and $Q_{n}^{right}(\tilde{\theta})$. It continues recursively until a stopping criterion is met, such as reaching the minimum number of samples at a node or when no further improvement in gain is possible. Mortality rate predictions are then made at these nodes with no further splits, called leaf nodes or terminal nodes. If $n$ is a terminal node, then the prediction for each county $i$ which reached this node is
@@
To evaluate the importance of each SVI component in the model, XGBoost utilizes the gain from eq. 1. This metric reflects the improvement in the model’s predictive performance when an SVI feature is used to split a decision node. At each node, the optimal split was chosen via eq. 2, which can be rewritten as follows:
@@
Here $J$ is the number of nodes which used feature $j$ to split. The features that contribute the most to improving the model’s predictive accuracy are those with the highest total gain. Consequently, SVI components with higher total gain are considered more important, as they most consistently reduce the prediction error across the model.
Autoencoder model
In this study, the autoencoder model was used to predict opioid-related mortality rates from 2011 to 2022 based on the previous year’s SVI data. The input data is first passed through the initial 1D convolutional layer, which aggregates the SVI variables before being fed into the encoder. The encoder compresses the data into a lower-dimensional representation, which is then passed to the decoder to produce the final vector containing yearly mortality rate predictions for each county in the nation.
To preserve the natural temporal structure of the opioid crisis, the model is trained sequentially on SVI data from 2010 to 2020, with 2015 held out as the validation set, and 2021 used for testing. To prevent overfitting, the model is allowed to train for up to 100 epochs, with early stopping triggered after 10 consecutive worse predictions on the validation set. The training process utilizes an $L^{1}$ loss function and a cyclical learning rate.
To describe the autoencoder model mathematically, again let $t\in\{2010,2011,\dots,2021\}$ denote a fixed year, and $X^{(t)}$ denote the input SVI data from year $t$. Let $\hat{Y}^{(t+1)}$ be the output vector of mortality rate predictions for year $t+1$, and let $A$ be the full neural network model. It consists of five layers: the initial 1D convolutional layer $C$, two encoding layers $E_{1}$ and $E_{2}$, and two decoding layers $D_{3}$ and $D_{4}$. Then the model can be expressed as follows:
@@
The convolutional layer $C$ learns an effective way to convolve the thirteen SVI components, producing a single vector of size 3144. This vector represents the aggregate state of the nation with respect to the SVI variables. The convolutional layer can be written as:

### v3

@@ 2.3  County-level characteristic data
2.4  XGBoost model
In this study, XGBoost was used to predict opioid-related mortality rates from 2011 to 2022 based on the previous year’s data. To use XGBoost, several parameters need to be selected, such as the number and depth of trees. To select these parameters, we employed a grid search algorithm with cross-validation, optimizing for the highest predictive accuracy. For training and testing, we used 5-fold cross-validation to randomly split the input data into five folds. In each fold, XGBoost trained on four subsets and predicted on one. This process was repeated across all five folds to generate predictions for every county, ensuring that the model was both trained and tested on the entire dataset.
To describe this model mathematically, fix a year $t\in\{2010,2011,\dots,2021\}$ and let $X^{(t)}=\big{[}x_{i,j}^{(t)}\big{]}\in\mathbb{R}^{3144\times 13}$ be the matrix of county-level characteristics input to the model in year $t$. Each row of matrix $X^{(t)}$ corresponds to a county, while each column corresponds to a characteristic. In particular, let $x_{i,*}^{(t)}$ denote the row vector containing all the feature data for county $i$, and let $x_{*,j}^{(t)}$ denote the column vector containing all county rates for variable $j$. Letting $Y^{(t+1)}=\big{(}y_{i}^{(t+1)}\big{)}\in\mathbb{R}^{3144}$ be the target vector of mortality rates in year $t+1$, we can define the full dataset $X^{(t)}\times Y^{(t+1)}$ passing through the model to be the set of all counties’ feature vectors together with their corresponding target values:
@@ 2.4  XGBoost model
Starting with the root node, an individual decision tree grows by recursively applying a splitting process to each node $n$ in the tree. Let the data at node $n$ be denoted by $Q_{n}$ with $s_{n}$ samples. Then $s_{n}$ is the number of counties which reached that node, and $Q_{n}\subseteq X^{(t)}\times Y^{(t+1)}$ is the subset of data for those counties. At each node, a potential split $\theta=(j,h_{n})$, consisting of a feature $j$ and a threshold $h_{n}$, partition $Q_{n}$ into two subsets:
$Q_{n}^{left}(\theta)$ $=\{(x_{i,*}^{(t)},y_{i}^{(t+1)})\,|\,x_{i,j}^{(t)}\leq h_{n}\},$
@@
[…]
@@
To evaluate the importance of each county-level characteristic in the model, XGBoost utilizes the gain from eq. 1. This metric reflects the improvement in the model’s predictive performance when a county-level characteristic is used to split a decision node. At each node, the optimal split was chosen via eq. 2, which can be rewritten as follows:
@@
Here $J$ is the number of nodes which used feature $j$ to split. The features that contribute the most to improving the model’s predictive accuracy are those with the highest total gain. Consequently, variables with higher total gain are considered more important, as they most consistently reduce the prediction error across the model.
2.5  Autoencoder model
In this study, the autoencoder model was used to predict opioid-related mortality rates from 2011 to 2022 based on the previous year’s data. The input data is first passed through the initial 1D convolutional layer, which aggregates the thirteen characteristics before being fed into the encoder. The encoder compresses the data into a lower-dimensional representation, which is then passed to the decoder to produce the final vector containing yearly mortality rate predictions for each county in the nation.
To preserve the natural temporal structure of the opioid crisis, the model is trained sequentially on data from 2010 to 2020, with 2015 held out as the validation set, and 2021 used for testing. To prevent overfitting, the model is allowed to train for up to 100 epochs, with early stopping triggered after 10 consecutive worse predictions on the validation set. The training process utilizes an $L^{1}$ loss function and a cyclical learning rate.
To describe the autoencoder model mathematically, again let $t\in\{2010,2011,\dots,2021\}$ denote a fixed year, and $X^{(t)}$ denote the input data from year $t$. Let $\hat{Y}^{(t+1)}$ be the output vector of mortality rate predictions for year $t+1$, and let $A$ be the full neural network model. It consists of five layers: the initial 1D convolutional layer $C$, two encoding layers $E_{1}$ and $E_{2}$, and two decoding layers $D_{3}$ and $D_{4}$. Then the model can be expressed as follows:
@@ 2.5  Autoencoder model
The convolutional layer $C$ learns an effective way to convolve the thirteen county-level characteristics, producing a single vector of size 3144. This vector represents the aggregate state of the nation with respect to the variables. The convolutional layer can be written as:
@@
3  Results
3.1  Data analysis
We sourced mortality rates from the Centers for Disease Control and Prevention’s Wide-ranging ONline Data for Epidemiologic Research (CDC WONDER) system [53], which suppresses death counts in counties with fewer than 10 deaths. Consequently, mortality rates were unavailable for nearly half of the nation’s counties in each year of the study period. To address this data censoring, we developed a novel methodology to impute missing values based on data from neighboring counties, as detailed in section 2.2.
Using this imputed dataset, our study begins by analyzing how the rates of each SVI component manifests in counties with anomalous mortality rates. For each year in the study period, we define anomalies as rates that fall within specific percentiles of a statistical distribution fitted to the nonzero mortality rates. Directly including the zero values in the fitting process poses significant challenges. For example, zero values account for more than 30% of the data in several years, creating a large spike at zero that distorts both parametric and empirical distributional fits.
To address these challenges, we analyzed counties with zero mortality rates separately from those with nonzero rates. We found that the characteristic patterns in the zero-rate counties and the anomalously low-rate counties were nearly identical. These findings will be discussed in greater detail later in this section, but we note here that these nearly indistinguishable patterns support our decision to focus on the distribution of nonzero rates. This approach allows us to define anomalous mortality while still capturing the broader social patterns present in the zero-rate counties.
Consequently, to define anomalous mortality rates, we restrict the set of candidate distributions to those defined on the domain $(0,\infty)$. To determine the best-fitting distribution, we evaluated Gamma, Weibull, inverse Gaussian, and lognormal distributions. The goodness-of-fit was assessed using the Kolmogorov-Smirnov test, Akaike Information Criterion, and Bayesian Information Criterion. Based on these statistical measures, the lognormal distribution consistently outperformed the other candidates, scoring as the best-fitting distribution each and every year in the study period. Using the fitted lognormal distributions, anomalies are defined annually in two categories as follows: “Hot anomalies” are counties with nonzero rates above the 98th percentile of the yearly fitted distribution, while “cold anomalies” are those whose nonzero rates fall below the 2nd percentile. To select the size of these tails, we assessed 1%, 2%, and 3% thresholds. In some years, a 1% tail produced no cold anomalies, whereas the 2% tail always provided them. While a 3% tail also could have been used, to maximize the anomalous nature of the rates, we opted for the tightest bounds that still yielded cold anomalies for analysis each year. Figure 3 shows how the anomalies in 2016 change with 1%, 2%, and 3% tails.
Figure 3: Impact of threshold selection on 2016 anomalies. (A) 2016 anomaly map using 1% tails. (B) 2016 anomaly map using 2% tails. (C) 2016 anomaly map using 3% tails. A 1% threshold produced no cold anomalies, while a 2% threshold provided the smallest bound that yielded cold anomalies. Increasing the threshold to 3% further raised the number of anomalies in both categories. Figures are best viewed in color.
To compare the rates of each characteristic within the anomalous counties, we rank the variables based on their mean rates. Specifically, we first calculate the mean rate within each year and then average these yearly means over the entire study period. Features are ordered from highest to lowest average mean rate in hot counties and from lowest to highest in cold counties. These rankings identify the variables most strongly correlated with anomalous mortality rates. The results are presented in figure 4, where yearly mean rates for each feature are shown using colored bars, and black bars indicate each variable’s average mean rate across all years of the study.
Figure 4: Mean rates of the county-level characteristics in the hot and cold anomalies. (A) Hot anomaly summary, where features are ranked from highest average mean rate down to lowest. (B) Cold anomaly summary, where features are ranked from lowest average mean rate down to highest. In both plots, each feature’s yearly mean rates are displayed using colored bars, while the black bars showcase each feature’s average mean rate across all years in the study. Figures are best viewed in color.
Figure 4a ranks the characteristics based on their average mean rates in the hot counties. Among these hot counties, the variable with the highest average mean rate is no vehicle (75.40), followed by below poverty (75.22), no high school diploma (73.48), unemployment (70.18), and mobile homes (68.61). The high ranking of these features means they tend to manifest their highest rates in counties with anomalously high opioid-related mortality. This suggests that opioid-related mortality may be exacerbated by high levels of these factors. Beyond these top five, rates noticeably decline, starting with single-parent households (54.61) and continuing to drop further down the list. At the bottom are limited English ability and minority status, meaning they have the lowest average rates in the hot counties. This suggests that these characteristics may contribute the least to the exceptionally high mortality rates observed in the hot counties.
Figure 4b ranks the variables based on their average mean rates in the cold counties. Among these cold counties, the characteristic with the lowest average mean rate is unemployment (35.58), followed by multi-unit structures (39.66), single-parent households (39.89), limited English ability (42.22), and no vehicle (42.84). The high ranking of these features means they tend to manifest their lowest rates in counties with anomalously low opioid-related mortality. This suggests that opioid-related mortality may be alleviated by lower levels of these factors. Beyond these top five, mean rates only gradually increase, except at the bottom, where a significant jump is observed for the variable aged 65 or older. This factor stands out with the highest average mean rate (61.74) in the cold anomalies, indicating that counties with anomalously low opioid-related mortality tend to have a significantly higher population of persons aged 65 or older.
As noted at the beginning of this section, the pattern observed in the zero mortality rate counties and the one just described for the cold counties is nearly indistinguishable. This is illustrated in fig. 5, where the only difference is a swap in order between no high school diploma and below poverty, both of which have average means that were nearly the same.
Figure 5: Mean rates of the county-level characteristics in the zero mortality and cold counties. (A) Zero mortality counties summary, where features are ranked from lowest average mean rate down to highest. (B) Cold counties summary, where features are also ranked from lowest average mean rate down to highest. In both plots, each feature’s yearly mean rates are displayed using colored bars, while the black bars showcase each feature’s average mean rate across all years in the study. Figures are best viewed in color.
In this data analysis, notably, only two variables—unemployment and lack of vehicle access—exhibited bidirectional relationships, with their highest rates occurring in counties with anomalously high mortality and their lowest rates in counties with anomalously low mortality. The fact that these features are observed at both extreme ends of the mortality rate spectrum suggests that they may play a particularly significant role: High levels may contribute to increased opioid-related mortality, while low levels may help reduce it. These empirical results provide a real-world foundation for the subsequent machine learning analysis, identifying which features are the key contributors to opioid-related mortality predictions.
3.2  Efficacy of machine learning models
This study employed two machine learning models to predict opioid-related mortality rates from 2011 to 2022 using the previous year’s data. XGBoost generates predictions for each county by combining the outputs from an ensemble of decision trees, where each tree iteratively corrects the errors of its predecessors through gradient-based optimization. The autoencoder model, on the other hand, processes input data through an initial 1D convolutional layer that aggregates the county-level characteristics before passing the data to an encoder, which compresses it into a lower-dimensional representation. This representation is then decoded to produce the yearly mortality rate predictions.
The efficacy of our machine learning models in forecasting county-level mortality rates is evaluated using both accuracy and error metrics. For each county, the error is measured as the absolute residual. To assess overall accuracy, the maximum absolute residual, which indicates the largest annual prediction error, is used as a benchmark: Each county’s error is normalized by this highest observed error across the country. The performance of the models is summarized in table 3, which presents the average errors, maximum errors, and average accuracy of both models for each predictive year in the study period; table 4 provides summary statistics for the mortality rate data to contextualize these results.
Year Avg Error Max Error Avg Accuracy
XGB AE XGB AE XGB AE
2011 5.79 4.07 100.18 72.41 94.22% 94.39%
2012 5.67 4.08 72.14 72.88 92.14% 94.40%
2013 5.86 3.97 78.39 55.25 92.52% 92.82%
2014 6.43 3.56 87.57 59.17 92.66% 93.98%
2015 6.65 3.26 112.40 68.05 94.09% 95.22%
2016 7.83 4.66 107.68 100.36 92.73% 95.36%
2017 8.34 4.15 126.24 72.20 93.39% 94.25%
2018 8.18 3.94 98.80 48.67 91.72% 91.91%
2019 8.43 4.02 97.56 60.61 91.36% 93.36%
2020 10.78 7.35 136.70 92.04 92.11% 92.02%
2021 13.84 11.60 161.62 184.12 92.06% 93.70%
2022 12.44 12.37 236.83 230.10 94.75% 94.62%
Table 3: Efficacy metric summary for both the XGBoost (XGB) and autoencoder (AE) models. For each county, the error is measured as the absolute residual, while the accuracy is calculated as the county’s error normalized by the highest observed error nationwide.
Year Mean Std Dev Min Q1 Median Q3 Max
2010 9.75 11.88 0.0 0.00 7.53 14.79 126.85
2011 10.33 11.87 0.0 0.00 8.66 15.48 132.37
2012 10.04 10.80 0.0 0.00 8.44 15.58 98.48
2013 10.45 10.55 0.0 0.00 9.07 16.26 100.59
2014 11.66 12.09 0.0 0.00 9.69 17.98 100.56
2015 12.43 12.46 0.0 0.00 10.72 19.28 141.16
2016 14.64 14.61 0.0 0.00 12.46 22.50 134.08
2017 15.75 15.72 0.0 0.00 13.54 23.92 152.70
2018 14.88 14.56 0.0 0.00 12.82 23.77 123.36
2019 15.37 14.85 0.0 0.52 13.12 24.05 119.13
2020 21.84 20.29 0.0 4.22 18.94 32.82 171.44
2021 27.29 23.93 0.0 8.80 24.32 39.30 212.38
2022 27.27 23.20 0.0 9.62 25.46 39.11 238.27
Table 4: Summary statistics for the mortality rates from 2010 to 2022.
While each year is treated as an independent test year by XGBoost, 2022 serves as the sole testing year for the autoencoder. This approach was chosen to preserve the natural temporal trajectory of the opioid crisis by training the autoencoder sequentially on data from 2010 to 2020, with 2021 as the testing year and predictions made for 2022. Accordingly, we present visual representations of the 2022 efficacy results using error histograms (fig. 6) and accuracy maps (fig. 7).
Figure 6: Error histograms for 2022 mortality rate predictions from both models. (A) Histogram displaying the distribution of XGBoosts’s absolute errors for the 2022 mortality rate predictions. (B) Histogram displaying the distribution of the autoencoder’s absolute errors for the 2022 mortality rate predictions. In both plots, the error stemming from Menominee County, Wisconsin, is indicated by the red arrow, while the error stemming from Bath County, Kentucky, is indicated by the orange arrow. Figures are best viewed in color.
Figure 7: Accuracy maps for 2022 mortality rate predictions from both models. (A) Accuracy map for XGBoost’s 2022 mortality rate predictions. (B) Accuracy map for the autoencoder’s 2022 mortality rate predictions. Both maps are color-coded to represent increasing intervals of accuracy by 5%, starting from dark red for the least accurate predictions, and progressing to dark green for the most accurate predictions. In both plots, Menominee County, Wisconsin, which had the largest 2022 error for both models, is shaded in the darkest red. To make it easier to locate, this county is also circled in the same color. Bath County, Kentucky, is shaded in orange on the autoencoder’s accuracy map and in yellow on XGBoost’s map, highlighting that the autoencoder had a higher error here than XGBoost. This county has also been circled in its respective color for easier visual identification. Figures are best viewed in color.
3.3  XGBoost feature importance
We used XGBoost to predict opioid-related mortality rates from 2011 to 2022 based on the previous year’s data. To assess the importance of each county-level characteristic, XGboost utilizes information gain:
This metric, detailed in section 2.4, quantifies the improvement in the model’s predictive performance when a feature is used to split a node in a decision tree. As a result, the features with the highest gain scores are those that most significantly improve the model’s ability to predict opioid-related mortality rates.
Figure 8: Feature importance rankings for both models. (A) XGBoost feature importance summary. (B) Autoencoder feature importance summary. Both plots display each feature’s yearly importance scores with colored bars, while black bars represent the average importance scores across all years. Features are ranked by their average importance, from highest to lowest in both plots. Figures are best viewed in color.
3.4  Autoencoder feature importance
In this study, the autoencoder model was used to predict opioid-related mortality rates from 2011 to 2022 based on the previous year’s data. To determine which county-level characteristics were most important in the autoencoder’s predictions, we applied a Shapley Gradient Explainer to the initial convolutional layer. The Shapley Gradient Explainer generates SHAP (Shapley Additive Explanations) values, which quantify the contribution of each feature to a model’s predictions [39, 40, 41]. By measuring how sensitive the autoencoder’s mortality predictions are to small changes in each feature, the SHAP values identify the most important components as those for which minor changes lead to the greatest impact on predictions.
The variable with the highest average SHAP value is crowding (0.00143), followed closely by no vehicle (0.00137) and multi-unit structures (0.00136). In fourth place is aged 65 or older (0.00131), from which SHAP values gradually decrease down to unemployment (0.00116) in seventh place. Beyond unemployment, there is a noticeable drop in the SHAP values for mobile homes and minority status. After which, there is another steep decline in importance moving down to the the last four characteristics. This reflects the much smaller contribution these four factors make to the model’s predictions. Notably, the high importance score of no vehicle corroborates the empirical findings from the data analysis.
Lastly, we note that SHAP values are relative to the model, data, and problem context; they reflect the contributions of each feature to the predictions compared to the baseline (i.e., the average prediction). Ranking these values within the model provides a reliable and interpretable measure of feature importance, offering meaningful insights into the influence of each variable on this specific model and dataset.
4  Discussion
Utilizing thirteen county-level characteristics, this study sought to identify which variables are most strongly correlated with opioid-related mortality. The data analysis portion of this study examined how the rates of each component manifested in counties with anomalous mortality. Notably only two factors—unemployment and lack of vehicle access—exhibited bidirectional relationships, with their highest rates observed in counties with anomalously high mortality and their lowest rates in counties with anomalously low mortality. This pattern suggests that reducing unemployment and improving vehicle access may lead to measurable reductions in opioid-related mortality, particularly in the most severely affected regions.
These empirical results provided real-world grounding for the subsequent machine learning analysis, which identified the characteristics that are key contributors to opioid-related mortality predictions. This portion of the study employed two machine learning models: XGBoost and an autoencoder. Feature importance was assessed using information gain for XGBoost and a Shapley gradient explainer for the autoencoder. While the feature importance rankings differed between the two models, this difference highlights the value of employing multiple approaches, as it revealed critical insights that might have been overlooked if only one model had been used.
Public health interventions could mitigate these impacts by prioritizing strategies that address both economic stability and transportation access. Workforce development programs that focus on providing stable, meaningful employment opportunities could play a critical role in reducing key stressors that contribute to opioid misuse, particularly in regions with high opioid-related mortality. Additionally, improving access to reliable and affordable transportation, especially in underserved communities and rural regions, could help alleviate the barriers faced by those without personal vehicles. By addressing these underlying structural challenges, targeted interventions could help reduce the impact of these two critical factors and may even promote long-term recovery in the nation’s most vulnerable areas.
This study has important limitations to note. Primarily, the missing mortality rates in the CDC WONDER dataset. Although we employed methods to address this, no approach can fully overcome the challenges posed by suppressed data. This limitation was sometimes evident when real data, missing in the previous year, became available the following year at a much higher rate than the previously imputed value. Such sharp increases occasionally led to large predictive errors in both models. Future research could apply the methodologies outlined in this study to a dataset with fewer missing values, if such data is available.
Another limitation is that the analysis is conducted at the county level. Using aggregate data may obscure important local differences, a phenomenon often referred to as the ecological fallacy. Additionally, this study does not stratify counties by urban or rural designation, but the same county-level characteristic can have markedly different implications in these two different contexts. For example, in urban areas such as New York City, low vehicle ownership rates may be a result of robust public transportation systems. Whereas in rural regions, such as Appalachia, lacking a vehicle can mean extreme isolation and limited access to healthcare. Future studies could explore such urban–rural dynamics to develop context-specific strategies for addressing transportation barriers.
Finally, while the thirteen SVI variables used in this study offer broad coverage of social factors, they do not capture all dimensions relevant to opioid-related mortality, such as: healthcare provider density, mental health service availability, incarceration rates, industry composition, long-term unemployment, college education rate, walkability, park access, or income inequality. For many of these factors, nationwide county-level data are either unavailable or limited to select years. These limitations render them impractical for the present design, as the timespan must align with that of the study period. Nevertheless, their omission constrains the model’s explanatory reach. If data availability permits, expanding the variable set in future work could help capture additional pathways and regional nuances in the opioid crisis.
5  Conclusion
Despite its limitations, this study provides valuable insight into some key structural factors driving the opioid crisis. By integrating exploratory data analysis with machine learning models, we identified vulnerabilities that may exacerbate opioid-related mortality when present at high levels, yet mitigate it when present at low levels. Across both methods of analysis, unemployment and lack of vehicle access consistently emerged as critical variables. These findings underscore the importance of addressing economic precarity and transportation barriers, particularly in regions most heavily affected by the epidemic. Additionally, while the limitations discussed above remain, the methods and results presented here offer a robust foundation for future research to build upon.